OpenStudy (anonymous):

Diameter of a cylindrical vessel is 60cm. It is filled with water such that a sphere of diameter 30cm immersed fully in it. Then, what is the increase in height of the water surface after putting the sphere in vessels

OpenStudy (anonymous):

is the cylinder fully filled or half full?

OpenStudy (anonymous):

please help me any1

OpenStudy (anonymous):

what is the height of the cylindrical vessel?

OpenStudy (anonymous):

only the above information is given NOTHING MORE IS GIVEN

OpenStudy (anonymous):

CAN U TELL ME M I RIGHT OR NOT I FOUND THE ANSWER BUT NOT SURE

OpenStudy (anonymous):

I FOUND : radius of the cylinder vessel is 30cm radius of the sphere is 15 cm volume of the sphere is 4/3 pi * (15)^3 cm^3 I SUPPOSED water level rises by h cm in the cylindrical vessel, Then the volume of cylinder of height h cm and radius 30cm= pi*(30)^2 * h cm^3 HENCE, VOLUME OF THE WATER DISPLACED BY THE SPHERE IS EQUAL TO VOLUME OF THE SPHERE: pi*(30)^2 * h=4/3(pi) * (15^3) h = (4/3 pi *15*15*15) div (pi * 30* 30 ) which givees H = 5cm IS IT RIGHT ANY1 TELL ME

OpenStudy (anonymous):

You are correct.

OpenStudy (anonymous):

As long as the water level in the cylinder is enough to *fully* immerse the sphere, the water level will rise based on the volume of the sphere as applied to the cylinder. If it is *not* filled enough to fully immerse the sphere, you'd need to determine the volume of the sphere that was immersed and apply the same technique. But rest assured, your solution is correct.